# Codimension-Two Spiral Spin-Liquid in the Effective Honeycomb-Lattice Compound Cs3Fe2Cl9

### v1

@@ X.5 The $J_{123}+D_{z}+J_{5}$ model
Figures S13-S15 summarize the simulation results for the $J_{123}+{D_{z}}+J_{5}$ model that considers a FM perturbation of $J_{5}=-0.008$ meV. This strength of $J_{5}$ is found to best describe the experimental phase diagram on a $12\times 12\times 4$ supercell. By carefully analyzing the field and temperature dependence of the magnetic susceptibility (see Fig. S14) and specific heat (see Fig. S15), we obtain the phase diagram shown in Fig. S13(a). The magnetic structure factors shown in Figs. S13(b-d) reproduce the $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in zero field and $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$ at $H=5$ T. Therefore, we conclude that FM $J_{5}$ is the main perturbative term to the $J_{123}+{D_{z}}$ Hamiltonian, leading to a $J_{123}+{D_{z}}+J_{5}$ model that is employed in the main text.
@@
Figure S16: Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ in phase III for the $J_{123}+D_{z}+J_{5}$ model. Calculations were performed on a $15\times 15\times 4$ supercell using the classical Monte Carlo simulations. The strength of $J_{5}$ is slightly increased to $-0.01$ meV to compensate for the size effect so that the phase boundaries are similar to that shown in Fig. 4(f) of the main text. The temperature and magnetic field are set to 1 K and 4 T, respectively. Calculations were performed over $5\times 10^{5}$ measuring sweeps after $5\times 10^{5}$ thermalization sweeps. (c-h) Similar magnetic structure factors in phases IV, V, and VI. The calculation parameters are $T=1$ K and $H=5.5$ T in phase IV, $T=3.1$ K and $H=4.0$ T in phase V, and $T=5.0$ K and $H=4.0$ T in phase VI.
X.5.1 Spiral-type orders in phases III and IV
@@
Comparisons of the direction-resolved magnetic structure factors and real-space spin configurations unveil the magnetic structures in each of the field-induced phases. Figure S16 compares the $\langle S_{z}S_{z}\rangle$ and $\langle S_{x}S_{x}\rangle$ components of the magnetic structure factors calculated for phases III, IV, V, and VI through the classical Monte Carlo simulations. In phases III and IV, both the $\langle S_{z}S_{z}\rangle$ and $\langle S_{x}S_{x}\rangle$ components exhibit long-range order, suggesting a spiral-type magnetic order that involves both in-plane and out-of-plane spin components. The magnetic structure factor along the $z$ direction is about 10 times higher than that along the $x$ direction, which indicates that the spiral order in phases III and IV is elongated along the $c$ axis due to the uniaxial SIA. Compared to that in phase III, the relatively stronger $2\bm{q}_{\textrm{III}}$ reflections in phase IV suggests stronger squaring-up effects in a higher magnetic field. This field dependence of the $2\bm{q}_{\textrm{III}}$ reflections is consistent with the experimental observations shown in Fig. S6.
X.5.2 SDW orders in phases V and VI
@@
As shown in Fig. S16(f) and (h), the $\langle S_{x}S_{x}\rangle$ component of the magnetic structure factors in phases V and VI exhibit diffuse patterns, while sharp Bragg peaks are observed in the $\langle S_{z}S_{z}\rangle$ component. This observation indicates that magnetic moments in phases V and VI are ordered only along the $c$ axis, leading to sinusoidally modulated SDW orders in these two phases. We also note that comparison of the $\langle S_{z}S_{z}\rangle$ component in phases III (Fig. S16(a)), IV (Fig. S16(c)), and V (Fig. S16(e)) reveals a similar magnetic propagation vector $\bm{q}_{\textrm{III}}$ and its high harmonics $2\bm{q}_{\textrm{III}}$, which reproduces the experimental observations shown in Fig. S16.
XI Analytical expressions for the field-induced phases
@@
Figure S17: (a) Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ using the ansatz in Eqn. (S2) for phase III. Calculations were performed for a two-dimensional $30\times 30$ honeycomb superlattice through Fourier transform. (c-d) Magnetic structure factors using the ansatz in Eqn. (S2) for phases IV. (e) Magnetic structure factor $\langle S_{z}S_{z}\rangle$ using the ansatz in Eqn. (S3) for phase V. (d) Magnetic structure factor $\langle S_{z}S_{z}\rangle$ using the ansatz in Eqn. (S3) for phase VI.
@@
XII Magnetic order in the magnetization plateau phase
A few possible magnetic orders have been proposed for the 1/2-magnetization plateau phase Ishii et al. (2021). Through classical Monte Carlo simulations, the order in the 1/2-magnetization plateau phase of the $J_{123}+{D_{z}}+J_{5}$ model is presented in Fig. S18. Further neutron diffraction experiments in fields above 6 T will be required to verify this magnetic order.
Figure S18: Magnetic order in phase VII viewed along the $c$ axis as determined from classical Monte Carlo simulations for the $J_{123}+{D_{z}}+J_{5}$ model. The bottom left (top right) part depicts the spin configuration for two sublattices (one sublattice).
XIII Impacts of the $J_{5}$ perturbation on the spin dynamics
Figure S19 presents the calculated INS spectra and diffuse neutron scattering pattern in the $(h,k,0)$ plane for the $J_{123}+D_{z}+J_{5}$ model with a perturbative $J_{5}=-0.008$ meV. The INS spectra are calculated by the linear spin wave theory. The diffuse neutron scattering pattern is calculated by the SCGA method at $T=5$ K. The calculated results only exhibit marginal difference compared to those of the $J_{123}+D_{z}$ model as presented in the main text.
Figure S19: (a) Calculated INS spectra and (b) diffuse neutron scattering pattern in the $(h,k,0)$ plane for the $J_{123}+D_{z}+J_{5}$ model with $J_{5}=-0.008$ meV. The diffuse neutron scattering pattern is calculated by the SCGA method at $T=5$ K.

### v2

@@ XII Phase diagram for the $J_{1\textrm{-}5}$-$D_{z}$ model
Figures S8-S10 summarize the simulation results for the $J_{1\textrm{-}5}$-$D_{z}$ model. By carefully analyzing the field and temperature dependence of the magnetic susceptibility (see Fig. S9) and specific heat (see Fig. S10), we obtain the phase diagram shown in Fig. S8(a). The typical magnetic structure factors shown in Figs. S8(b-d) reproduce the $\bm{q}_{\textrm{I}}=(\frac{1}{2},0,0)$ in phase I and VII at $H=$ 0 and 10 T, respectively, and $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$ in phase IV at $H=5$ T.
Figure S8: (a) Theoretical phase diagram for the $J_{1\textrm{-}5}$-$D_{z}$ model calculated by the classical Monte Carlo simulations. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1.5$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T. Calculations of the magnetic structure factor were performed on a $16\times 16\times 4$ supercell with $2\times 10^{5}$ thermalization sweeps and $1\times 10^{5}$ measurement sweeps. White arrows indicate the positions of weak magnetic Bragg peaks.
Figure S11 compares the experimental and theoretical magnetization curves calculated by classical Monte Carlo simulations for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model. The identified critical fields are largely consistent with those in the phase diagram in Fig. S8(a). The existence of the 1/2 magnetization plateau phase is well reproduced in our simulations.
Figure S9: Theoretical magnetization curves calculated by the classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model. The strength of the applied field for each curve is indicated on the left. Data are successively shifted by 0.003 arb. units along the $y$ axis for clarity. The dashed line at $T=1.8$ K marks the lower limit of the measuring temperature in Ref. Ishii et al. (2021).
Figure S10: Theoretical heat capacity $C_{p}$ calculated by the classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model in the field regime of (a) [0, 3.2], (b) [4, 7], and (c) [8, 14] T. Data are normalized by the maximal $C_{p}$ value, $C_{p}^{\textrm{max}}$. The strength of the applied field for each curve is indicated on the left. In each panel, data are successively shifted by 1 unit along the $y$ axis for clarity. The dashed line at $T=2.0$ K marks the lower limit of the measuring temperature in Ref. Ishii et al. (2021).
Figure S11: Experimental (a) and theoretical (b) magnetization curves. The experimental data, together with the phase boundaries indicated by the dashed lines, are adapted from Ref. Ishii et al. (2021). The theoretical data were obtained from classical Monte Carlo simulations using the $J_{1\textrm{-}5}$-$D_{z}$ model.
@@
Figure S12: Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ in phase III for the $J_{1\textrm{-}5}$-$D_{z}$ model. Calculations were performed on a $16\times 16\times 4$ supercell using the classical Monte Carlo simulations. The temperature and magnetic field are set to 2.33 K and 4 T, respectively. Calculations were performed over $1\times 10^{5}$ measuring sweeps after $2\times 10^{5}$ thermalization sweeps. (c-h) Similar magnetic structure factors in phases IV, V, and VI. The calculation parameters are $T=2.33$ K and $H=5$ T in phase IV, $T=4.17$ K and $H=4.0$ T in phase V, and $T=5.58$ K and $H=5.0$ T in phase VI.
XII.1 Spiral-type orders in phases III and IV
@@
Comparisons of the direction-resolved magnetic structure factors and real-space spin configurations unveil the magnetic structures in each of the field-induced phases. Figure S12 compares the $\langle S_{z}S_{z}\rangle$ and $\langle S_{x}S_{x}\rangle$ components of the magnetic structure factors calculated for phases III, IV, V, and VI through the classical Monte Carlo simulations. In phases III and IV, both the $\langle S_{z}S_{z}\rangle$ and $\langle S_{x}S_{x}\rangle$ components exhibit long-range order, suggesting a spiral-type magnetic order that involves both in-plane and out-of-plane spin components. The structure factor along the $z$ direction is about 10 times higher than that along the $x$ direction, which indicates that the spiral order in phases III and IV is elongated along the $c$ axis due to the uniaxial SIA. Compared to that in phase III, the relatively stronger $2\bm{q}_{\textrm{III}}$ reflections in phase IV suggest stronger squaring-up effects in a higher magnetic field. This field dependence of the $2\bm{q}_{\textrm{III}}$ reflections is consistent with the experimental observations shown in Fig. S3.
XII.2 SDW orders in phases V and VI
@@
As shown in Fig. S12(f) and (h), the $\langle S_{x}S_{x}\rangle$ component of the magnetic structure factors in phases V and VI exhibits diffuse patterns, while sharp Bragg peaks are observed in the $\langle S_{z}S_{z}\rangle$ component. This observation indicates that magnetic moments in phases V and VI are ordered only along the $c$ axis, leading to sinusoidally modulated SDW orders in these two phases. We also note that comparison of the $\langle S_{z}S_{z}\rangle$ component in phases III (Fig. S12(a)), IV (Fig. S12(c)), and V (Fig. S12(e)) reveals a similar magnetic propagation vector $\bm{q}_{\textrm{III}}$ and its high harmonics $2\bm{q}_{\textrm{III}}$, which reproduces the experimental observations shown in Fig. S3.
XII.3 Magnetic orders in phases VII and VIII
A few possible magnetic orders have been proposed for the 1/2-magnetization plateau phase Ishii et al. (2021). Through classical Monte Carlo simulations, the order in the 1/2-magnetization plateau phase of the $J_{1\textrm{-}5}$-$D_{z}$ model is presented in Fig. S13(a). In the same figure, the predicted magnetic order in phase VIII of the $J_{1\textrm{-}5}$-$D_{z}$ model is presented in Fig. S13(b). Further neutron diffraction experiments in fields above 6 T will be required to verify these magnetic orders.
Figure S13: Magnetic orders in phases (a) VII, (b) VIII, and (c) I’ viewed along the $c$ axis as determined from classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model. The bottom left (top right) part depicts the spin configuration for two sublattices (one sublattice).
XII.4 Additional intermediate phases
As compared in Fig. 4 of the main text, the calculated phase diagram contains two intermediate phases that are possibly not observed in the experiment. In the regime of the experimental phase II, the calculated structure is a double-$\mathbf{q}$ phase that is denoted as I’, where the ordering vectors are the same as those in phase I. The simulated magnetic order in this phase is presented in Fig. S13(c). Close to the lower phase boundary of the 1/2-plateau phase VII, another transitional phase, designated VI’, is identified in the calculated phase diagram. The magnetic order in phase VI’ is largely the same as that of phase VI, while the $\langle S_{x}S_{x}\rangle$ and $\langle S_{y}S_{y}\rangle$ components is enhanced in phase VI’ due to its relatively low temperatures.
In our classical Monte Carlo simulations, we did not manage to remove these additional phases, especially phase VI’, by introducing further perturbations to the $J_{1\textrm{-}5}$-$D_{z}$ model. Therefore, we propose that the absence of this transitional phase in experiments might be due to quantum fluctuation that is not considered in our classical Monte Carlo calculations. On the antiferromagnetic triangular lattice model, quantum fluctuation is known to stabilizes a 1/3 magnetization plateau with a collinear up-up-down magnetic order even in the case of $S=5/2$ Coletta et al. (2016). Similarly, in Cs3Fe2Cl9, quantum fluctuation may enhance the stability of the 1/2 magnetization plateau phase and thus remove the transitional VI’ phase obtained in classical Monte Carlo simulations.
XIII Analytical expressions for the field-induced phases
@@
Figure S14: (a) Theoretical magnetic structure factors (a) $\langle S_{z}S_{z}\rangle$ and (b) $\langle S_{x}S_{x}\rangle$ using the ansatz in Eqn. (S2) for phase III. Calculations were performed for a two-dimensional $30\times 30$ honeycomb superlattice through the Fourier transform. (c-d) Magnetic structure factors using the ansatz in Eqn. (S2) for phases IV. (e) Magnetic structure factor $\langle S_{z}S_{z}\rangle$ using the ansatz in Eqn. (S3) for phase V. (d) Magnetic structure factor $\langle S_{z}S_{z}\rangle$ using the ansatz in Eqn. (S3) for phase VI.
@@
Figure S15: (a) Calculated diffuse scattering pattern in the (h,k,0h,~k,~0) plane using the fitted $J_{123}$-$D_{z}$ model at $T=6$ K. (b) Temperature evolution of the calculated spin correlation function $\langle\bm{S}\cdot\bm{S}\rangle$ along ($h$, $h$, 0) around (0.5, 0.5, 0). The integration area is outlined in panel (a) by a dashed rectangle. Black dots indicate the peak positions at the corresponding temperatures through Gaussian fits. (c) Temperature evolution of the calculated spin correlations $\langle S_{x}\cdot S_{x}\rangle$ (red triangles), $\langle S_{z}\cdot S_{z}\rangle$ (yellow circles), and their ratio $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ (black dots). The integration area is outlined in panel (a) by a dashed-line hexagon. Red line over the $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ data is the fitting curve to the Arrhenius law $A_{1}\exp(-\frac{T-T_{N}}{\Delta_{\textrm{SIA}}})+A_{2}$ with fitted parameters of $A_{1}=-0.561(4)$, $\Delta_{\textrm{SIA}}=3.93(7)$ K, and $A_{2}=0.913(2)$.
XIV Effects of the uniaxial anisotropy on the SSL
Classical Monte Carlo simulations for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model were performed to study the effects of the SIA on the SSL in Cs3Fe2Cl9. As a reference, the diffuse scattering pattern at $T=6$ K calculated by the SCGA method is shown in Fig. S15. Through classical Monte Carlo simulations that incorporate critical scattering near $T_{N}$, Fig. S15(b) summarizes the temperature evolution of the calculated scattering intensity along ($h+\frac{1}{2}$, $h+\frac{1}{2}$, 0), where the integration range is outlined by dashed rectangle in Fig. S15(b). For the fitted $J_{1\textrm{-}5}$-$D_{z}$ model, $T_{N}$ is found to be $\sim 5.75$ K. The separated peaks at $T/T_{N}>1$ correspond to thermally stabilized spiral correlations in the SSL state, while their gradually reduced separation with decreasing $T$ reveals the second-order character of the transition into the LRO. Figure S15(c) compares the in-plane ($x$) and out-of-plane ($z$) components of the calculated spin correlations as a function of $T$. Due to the existence of uniaxial SIA, the evolution of $\langle S_{x}\cdot S_{x}\rangle$ is nonmonotonic, and the ratio between the in-plane and out-of-plane components exhibits an Arrhenius-like behavior with an activation gap of $\Delta_{\textrm{SIA}}=3.93(7)$ K in a wide temperature regime. The value of the activation energy is close to the estimation from the single ion anisotropic energy of $D_{z}S(S+1)\sim 3.27$ K that is required to flip a single spin from the easy axis along $c$ to the $ab$ plane.